OpenStudy (anonymous):

A carpenter is building a rectangular room with a fixed perimeter of 140 ft. What dimensions would yeild the maximum area? what is the maximum area?

OpenStudy (anonymous):

rectangle length L, width W perimeter = 2L + 2W = 140 so L = (140-2W)/2 so area A =W(140-2W)/2 = 70W - W^2 to find max area, differentiate A wrt W, set to zero and solve